OpenStudy (anonymous):

What is the simplified form of (30x^6)/(14y^5) * (7y^2)/6x^4)

OpenStudy (anonymous):

\[\left(\frac{15 x^6}{7 y^5}\right)*\left(\frac{7 y^2}{6}\right)*x^4=\frac{5 x^6}{2 y^3}*x^4=\frac{5 x^{10}}{2 y^3} \]